A social psychologist hypothesizes that the more initial attraction a person has to another predicts more anxiety before their first date. Below are the first date data from a sample of participants. What can the psychologist conclude with an α of 0.01?

Attraction Anxiety
2
6
1
3
6
9
6
6
3
8
11
5
8
10
11
8
8
7


a) What is the appropriate statistic?
Slope

Chi-Square

Compute the statistic selected above:
Answer #1

Appropriate statistic- slope

X Y XY
2 8 16 4 64
6 11 66 36 121
1 5 5 1 25
3 8 24 9 64
6 10 60 36 100
9 11 99 81 121
6 8 48 36 64
6 8 48 36 64
3 7 21 9 49
Ʃx = Ʃy = Ʃxy = Ʃx² = Ʃy² =
42 76 387 248 672
Sample size, n = 9
x̅ = Ʃx/n = 42/9 = 4.66666667
y̅ = Ʃy/n = 76/9 = 8.44444444
SSxx = Ʃx² - (Ʃx)²/n = 248 - (42)²/9 = 52
SSyy = Ʃy² - (Ʃy)²/n = 672 - (76)²/9 = 30.2222222
SSxy = Ʃxy - (Ʃx)(Ʃy)/n = 387 - (42)(76)/9 = 32.3333333

Slope, b = SSxy/SSxx = 32.33333/52 = 0.6217949

y-intercept, a = y̅ -b* x̅ = 8.44444 - (0.62179)*4.66667 = 5.542735

Regression equation :

ŷ = 5.5427 + (0.6218) x

Sum of Square error, SSE = SSyy -SSxy²/SSxx = 30.22222 - (32.33333)²/52 = 10.1175214

Standard error, se = √(SSE/(n-2)) = √(10.11752/(9-2)) = 1.20223

Standard error for slope, se(b1) = se/√SSxx = 1.20223/√52 = 0.16672

Slope Hypothesis test:

Null and alternative hypothesis:

Ho: β₁ = 0 ; Ha: β₁ > 0

Test statistic:

t = b/(seb1) = 3.7296

df = n-2 = 7

p-value = T.DIST.2T(ABS(3.7296), 7) = 0.0037

Conclusion:

p-value < α Reject the null hypothesis.

There is enough evidence to conclude that there is a positive relationship between the two variables.